Grade 11Algebra

The 4th term of GP is square of its second term and the first term is -3 determined the 7 term

Abhi
8 Years agoGrade 11

1 Answer

lakshya pokharna
8 Years ago
formula for n term of gp is ar^n-1so as given in question (ar^4-1)=(ar)^2 r=1a=-3=-3(1)^(7-1)=-3(1)=-3 here ^ is power or raise to power